OpenStudy (anonymous):

Given f(x) = –5x + 3, find f(a + 1).

OpenStudy (anx):

just plug (a+1) in for x f(a+1) = -5(a+1)+3 distribute = -5a -5 +3 = -5a -2

OpenStudy (anonymous):

Replace x by a+1 in -5x+3

OpenStudy (timmyk14):

Darn, you guys beat me to it!